OpenStudy (anonymous):

Solve. 6x + 5 2x -11=90

OpenStudy (anonymous):

6x + 5 + 2x - 11 =90

OpenStudy (alexwee123):

first add like terms... 6x + 5 + 2x - 11 =90 8x-6=90 what do you think the next step is @Rainbow.ninja

OpenStudy (anonymous):

6 + 60?

OpenStudy (alexwee123):

yes 8x=90+6 then what next?

OpenStudy (anonymous):

8/96?

OpenStudy (alexwee123):

you got it the other way around... so what does 96/8 equal?

OpenStudy (anonymous):

12?

OpenStudy (alexwee123):

yes good job so x=12

OpenStudy (anonymous):

Thanks~